# Philip ${}^{*\,\sharp}$ Yandex, MIPT &Aleksandr Beznosikov

### v1

@@ 3.2 Advantage of the Full-Rank Updates
The insight from the Section 3.1 suggests that the training of language models performs significantly better when the entire parameter space is utilized during the training process. Given the importance of updating parameters in all directions, this poses the question: Is it optimal to use low-rank updates, as employed by methods such as GaLore (Zhao et al., 2024a), ReLoRA (Lialin et al., 2023), and BAdam (Luo et al., 2024)? Using low-rank updates means the effective rank of the update is significantly smaller than the full dimensionality of the parameter space, inevitably leading to a loss of valuable information contained in the gradient.
@@
However, the method to leverage the full-rank gradient for updating parameters is not readily obvious. Using algorithms like Adam (Kingma, 2014) is not an option due to the memory overhead they introduce, which is precisely what we aim to avoid. An alternative approach is to use state-free optimizers such as SGD or signSGD (Bernstein et al., 2018). Unfortunately, SGD have been shown to be ineffective for training transformer models, as shown in Zhang et al. (2020); Pan & Li (2023).
@@
Nevertheless, a recent study Zhao et al. (2024b) suggests a promising methodology: while SGDM doesn’t generally work well with transformers, using SGDM for the majority of parameters and Adam for a selected subset can lead to effective training. This raises the question: could a hybrid approach using SGD or signSGD instead of SGDM be viable? If the key subset of parameters is handled by advanced algorithms, can the other parameters be trained effectively with state-free optimizers?
@@
To address this question, we conducted an experiment on LLaMA-130m, where we utilized the Adam (Kingma, 2014) for state-full parameters and signSGD (Bernstein et al., 2018) for state-free parameters. A detailed description of the experimental setup can be found at Section A.1. The result is highlighted as the entry “Full-rank Random” in Table 1. Full-rank updates significantly enhance performance, approaching the efficiency of the memory-intensive Adam optimizer, which serves as a upper bound in terms of performance. These findings underscore the potential of state-free algorithms for updating a substantial portion of the parameter space, paving the way for efficient, scalable optimization methods that deliver high performance without the significant memory costs traditionally associated with state-of-the-art optimizers.
3.3 FRUGAL as Optimization Framework
@@
The setup outlined at the conclusion of the Section 3.2 results in a general framework for memory-efficient optimization. Briefly, it operates as follows: the entire space is partitioned into state-full and state-free subspaces. The state-full subspace is updated using advanced algorithms, while the state-free subspace is updated using a state-free method. After a certain number of steps, the state-full subspace is changed to better explore the optimization space. However, determining the optimal state-free optimizer and the projection method onto the state-full subspace is not readily apparent. In this section, we strive to find the optimal configuration.
@@
State-free optimizer. We conducted a preliminary experiment updating all parameters using state-free algorithms to choose between SGD and signSGD (Bernstein et al., 2018). Table 8 presents these results. After testing various learning rates, we found that signSGD consistently outperforms SGD, leading us to favor signSGD. We attribute this performance to the similarities between signSGD and Adam (Kingma, 2014), as noted in Balles & Hennig (2018); Balles et al. (2020); Kunstner et al. (2023). Additionally, signSGD produces updates of similar magnitude to those generated by Adam, which simplifies the calibration of the learning rate for state-free parameters.
@@
Projection type. When selecting a projection method, it is crucial to strike a balance between quality and memory efficiency. For instance, the approach using the projection matrix from SVD decomposition, as in GaLore (Zhao et al., 2024a), better preserves the information embedded in the gradient but requires additional memory for storing projection matrices and computational resources for performing the SVD. To reduce computational demands, one could employ random coordinate projection denoted as RandK, but this requires additional memory or recomputation [^3] See Appendix B for discussion on the memory requirements for different projection methods.. A more structured alternative is to select not random matrix elements but entire random columns or rows. The most aggressive approach follows the method from BAdam, wherein an entire block is chosen as the state-full subspace.
The performance results obtained with all these variants are presented in Table 1. Full-rank SVD outperforms both Full-rank RandK and Full-rank Block methods, demonstrating comparable performance. The superior performance of SVD projections can be explained by their ability to extract the principal information from the gradient. Nonetheless, a downside is the increased compute and memory demand from SVD. Therefore, we opt for the blockwise selection, as it is the most memory-efficient — requiring only the storage of active block indices.
A formal description of the final algorithm is presented in Algorithm 1. It’s worth noting that this framework allows for variation not only in the State-Full algorithm but also in the choice of projection and State-Free optimizer. In our experiments in Section 5, we use a specific variant with Adam as the State-Full optimizer, signSGD as the State-Free optimizer, and blockwise projection.
@@
For Line 7, state projection, in Algorithm 1, we note that if the projection does not change, i.e., $P_{k,i}=P_{k-1,i}$, then $P_{k,i}(P^{-1}_{k-1,i}(s))=s$. Thus, we only need to project states when the projection changes from one round to another. However, our preliminary experiments with RandK selection showed that resetting states performs comparably to projection. Therefore, we could replace this projection with state resetting when the projection changes, which also aligns with blockwise subspace selection. However, either resetting or projecting states is important since we want projected gradients and optimizer states to reside in the same space. For instance, GaLore ignores this step, which leads to degraded performance when projections are updated frequently; see Appendix C for details.
@@
2:     Compute stochastic gradient $\tilde{g}^{k}\leftarrow\nabla f_{\zeta^{k}}(x^{k})$;
3:     Update momentum vector $\tilde{m}_{j}^{k}\leftarrow(1-\beta)\tilde{g}_{j}^{k}+\beta\begin{cases}\tilde{m}_{j}^{k-1}&\text{if }j\in J_{k},\\ 0&\text{otherwise;}\end{cases}$
4:     Compute update vector $\tilde{u}_{j}^{k}\leftarrow\begin{cases}\tilde{m}_{j}^{k}&\text{if }j\in J_{k},\\ \tilde{g}_{j}^{k}&\text{otherwise;}\end{cases}$
5:     Update iterate $x^{k+1}\leftarrow x^{k}-\alpha^{k}\tilde{u}^{k}$;
@@ 4 Theoretical Results
For the theoretical analysis, we consider the case where the State-Free optimizer is SGD and the State-Full optimizer is SGD with momentum (SGDM). For the projection, we use coordinate-wise projection. This special case of FRUGAL is provided in Algorithm 2. We minimize the following objective
$\textstyle\min_{x\in\mathbb{R}^{d}}\big{\{}f(x):=\mathbb{E}_{\zeta}[f_{\zeta}(x)]\big{\}},$ (1)
where we access $f$ via a stochastic oracle that takes $x$ as input and returns $(f_{\zeta}(x),\nabla f_{\zeta}(x))$.
@@ 4.1 Notation and Preliminaries
Assumption 1.
  1. 1.
    Smoothness: The objective $f(x)$ in equation 1 is $L$-smooth.
  2. 2.
  3. 3.
    Independent samples: The random samples $\{\zeta_{k}\}_{k=1}^{\infty}$ are independent.
  4. 4.
4.2 Convergence of Algorithm 2
Below, we present the main convergence theorem.
Theorem 1.
Let Assumption 1 hold and $\alpha^{k}=\alpha\leq\frac{1-\beta}{L(4-\beta+\beta^{2})}$. Then, the iterates of Algorithm 2 satisfy
$\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}[\|g^{i}\|^{2}]$ $=\mathcal{O}\left(\frac{f(x^{1})-f^{*}}{k\alpha}+L\alpha\sigma^{2}\left(1+\frac{\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\beta}{(1-\beta)}\right)\right),$ (2)
The proof is deferred to Appendix E. Let us analyze the obtained result. Firstly, if $J_{k}=[d]$ or $J_{k}=\emptyset$, Algorithm 2 becomes SGDM and SGD, respectively. In this case, we have $\bar{p}^{k}_{\min}=1$ for SGDM and $\hat{p}^{k}_{\max}=0$ for SGD. Therefore, the resulting rate is $\mathcal{O}\left(\nicefrac{{1}}{{k\alpha}}+L\alpha\sigma^{2}\right)$, which recovers the best-known rate for both SGD and SGDM under these assumptions Liu et al. (2020). Furthermore, if at each step each coordinate is sampled independently with probability $p$, we have $\bar{p}^{k}_{\min}=\hat{p}^{k}_{\max}=p$. Therefore, we recover the same rate if $p=\mathcal{O}\left(1-\beta\right)$ or $p=\mathcal{O}\left(\beta\right)$. Finally, in the worst case (e.g., $J_{k}$ is deterministic and $0<|J_{k}|<d$), we have $\bar{p}^{k}_{\min}=0$ and $\hat{p}^{k}_{\max}=1$. Thus, the rate becomes $\mathcal{O}\left(\nicefrac{{1}}{{k\alpha}}+\nicefrac{{L\alpha\sigma^{2}}}{{1-\beta}}\right)$, which is worse by a factor of $\nicefrac{{1}}{{1-\beta}}$. However, this is expected since the bias from momentum is not outweighed by the variance reduction effect, as only the coordinates with momentum enjoy reduced variance; see Lemmas 1 and 2 in the appendix for details.
5 Experiments
This section presents the main experimental results of the paper. To evaluate the performance of FRUGAL, we conducted experiments both on the pre-training and fine-tuning of language models.
5.1 Pre-training experiments
Setup. The core setup for pre-training is taken from the Zhao et al. (2024a). We utilize LLaMA-based (Touvron et al., 2023a) model architectures with up to 1B parameters and train them on the Colossal Clean Crawled Corpus (C4) dataset (Raffel et al., 2020). The C4 dataset is intended for pre-training, making this setup a good approximation of real-world applications. A detailed description of the setup can be found in Section A.1. However, we made several modifications that we would like to discuss in detail below.
  • Training Duration. The training approach in Zhao et al. (2024a) aligns with the empirical rule from scaling laws (Hoffmann et al., 2022), which suggests using approximately 20 times the model size in tokens for training. However, this number of tokens is far from achieving convergence. In practice, models are typically trained for significantly longer periods (Touvron et al., 2023b; Zhang et al., 2024b). One reason for this discrepancy is that the original scaling laws do not account for the inference of the model after training. Adjustments to scaling laws considering this parameter are discussed, for example, in (Sardana & Frankle, 2023). For our experiments we chose 200k steps for the 60M and 130M models, 240k for 350M model and 300k for the 1B model.
  • Learning Rate. The authors of GaLore suggested using different learning rates for fixed unprojectable parameters (Embeddings, RMSNorms (Zhang & Sennrich, 2019), Logits) and the remaining projectable parameters (attention and FFN weights modules weights). However, introducing additional hyperparameters complicates the use of the algorithm. Since both sets of parameters are state-full and trained using the same optimization algorithm, we always used the same learning rate for them in FRUGAL and BAdam. For GaLore learning rate see Section 5.1.
  • Mixed Precision instead of the pure bfloat16 training. Pure 16-bit training has been shown to potentially compromise model convergence and accuracy (Zamirai et al., 2020). This degradation stems from storing both the model weights and optimizer statistics in reduced precision formats such as float16 or bfloat16. However, these formats often lack sufficient precision in representing floating-point numbers. Consequently, mixed precision training has become a more common approach for training language models (Le Scao et al., 2023; Almazrouei et al., 2023)). While training in pure 16-bit format is also possible, stochastic rounding (Gupta et al., 2015; Zamirai et al., 2020) is often employed to mitigate the aforementioned issue. Given that the goal of this research is to identify the optimal optimization algorithm, we deemed it more appropriate to compare optimizers in a transparent and stable setup that does not require auxiliary tricks. Hence, we primarily used Mixed Precision training for its illustrative value in understanding each method’s potential. However, for completeness, we also conducted experiments in pure bfloat16 format, detailed in our ablation study Section 5.1.2.
@@ 5.1 Pre-training experiments
Table 2: Comparison of validation perplexity and memory estimation for various optimization methods across LLaMA model scales trained on C4. Lower $\downarrow$ is better. In parentheses, we also indicate the additional memory overhead introduced by the optimization algorithm. The values are calculated assuming that each float value occupies 4 bytes (float32). Note that Embeddings, RMSNorms, and Logits are always trained using state-full Adam.
@@
5.1.1 Comparison to existing memory-efficient algorithms
@@
To begin, we present the results of comparing FRUGAL with existing memory-efficient methods across four sizes of LLaMA-based architectures: 60M, 130M, 350M, and 1B [^4] See preliminary experimental results with LLaMA 7B in Appendix D.
Baselines. We use the following methods as baselines for our approach:
  • Full-rank Training. Training using memory-inefficient Adam. Weights, gradients, and statistics are stored and computed for all parameters. This serves as an upper bound for model performance.
  • BAdam. Luo et al. (2024) proposed a block coordinate descent (BCD)-type optimization method termed BAdam. The parameters are divided into blocks, which are then updated one by one using Adam. Similar to GaLore, the optimized block is updated every $T$ steps. Although this method was initially proposed only for fine-tuning, it is the closest method to our FRUGAL. Unlike BAdam, in our algorithm, state-free blocks are not frozen but are updated using signSGD.
  • Other Algorithms. Among other relevant methods, ReLoRA (Lialin et al., 2023) and MicroAdam (Modoranu et al., 2024) can also be highlighted. However, we did not include them for comparison in this paper for the following reasons: 1. ReLoRA was evaluated in (Zhao et al., 2024a), where it significantly underperformed compared to GaLore with the same memory budget. 2. MicroAdam. Its current implementation only supports bfloat16 master weights, whereas our main experiments conducted with mixed precision.
Main results. The results of our experiments are presented in Table 2, which includes both validation perplexity and memory footprint estimations for each method. We compared all memory-efficient methods under the same memory budget, specifically with a density $\rho=0.25$. Here, $\rho$ refers to the proportion of state-full parameters (see Section A.1 for a detailed explanation).
We conducted a grid search to determine the optimal learning rate for Adam, which we then applied uniformly to FRUGAL and BAdam (Luo et al., 2024). For GaLore (Zhao et al., 2024a), we found that using this same learning rate produced better results than the rate originally suggested in their paper. This discrepancy might be attributed to our experiments involving a significantly larger number of training steps than those for which GaLore’s original learning rate was optimized.
Table 2 demonstrates that FRUGAL significantly outperforms the memory-efficient baselines across all model sizes with the same memory budget, coming close to the performance of Adam.
@@
Table 3: Perplexity of LLaMA-130M models pre-trained on C4 for 100k iterations (10B tokens). The leftmost column indicates the modules moved to the state-free set and trained using signSGD. The results show that Logits, unlike Embeddings and RMSNorms, are exceptionally responsive to the choice of optimization algorithm from Adam to signSGD.
@@
Linear, Logits 34.66
Zero-density training. Table 2 also reveals a surprising result: FRUGAL with $\rho=0.0$ significantly outperforms both GaLore and BAdam, even when these competing methods use a higher density of $\rho=0.25$. Essentially, in this case, the parameters are divided into two parts — a state-full part consisting of the Embeddings, RMSNorms, and Logits, and a state-free part consisting of all other parameters. This division remains fixed throughout the training. We conducted additional experiments to determine the maximum subset of parameters that can be trained with a state-free optimizer without significant quality degradation. We systematically moved different combinations of the Embeddings, RMSNorms, and Logits from the state-full to the state-free set and observed the results during the training of LLaMA-130M. Table 3 reveals that the Logits demonstrates a dramatically higher sensitivity, with changes to its optimizer resulting in severe performance degradation. This finding aligns with results from (Zhao et al., 2024b), where the authors demonstrated that most parameters can be trained using SGDM, but the Logits require training with Adam.
5.1.2 Ablation study
We also conducted additional experiments to verify the robustness of our framework to various hyperparameters. Firstly, an ablation study on the state-full subspace update frequency $T$ in Table 10 shows that the performance keeps improving up to $T=200$. We note that, unlike in Zhao et al. (2024a), the perplexity does not significantly decrease even when reducing the update frequency to $T=10$ ($\sim 0.2$ drop vs. $\sim 4.$ drop for GaLore). A detailed explanation for this result can be found in Appendix C. Second, when using other schedulers, the performance gap between FRUGAL and baselines remains consistent, as shown in Tables 6 and 6. Then, the results of training in pure bfloat16 are presented in Table 8, demonstrating consistency with our main experiments in Table 2, i.e., FRUGAL significantly outperforms the baselines across these variations. We also conducted experiments to show how perplexity changes with varying $\rho$, and the results are presented in Table 11. Finally, we conducted an experiment to compare different strategies for selecting state-full blocks during training. The results in Table 10 show that there is no significant difference between random and structured block selection.
5.2 Fine-tuning experiments

### v2

@@ 3.2 Advantage of the Full-Rank Updates
The insight from Section 3.1 suggests that the training of language models performs significantly better when the entire parameter space is explored during the training process. Given the importance of updating parameters in all directions, this poses the question: Is it optimal to use low-rank updates, as employed by methods such as GaLore, ReLoRA, and BAdam? The effective rank of low-rank updates is significantly smaller than the full dimensionality of the parameter space, inevitably leading to a loss of valuable information contained in the gradient.
@@
However, the method to leverage the full-rank gradient for updating parameters is not readily obvious. Using algorithms like Adam (Kingma, 2014) is not an option due to the memory overhead they introduce, which is exactly what we aim to avoid. An alternative approach is to use state-free optimizers such as SGD or signSGD (Bernstein et al., 2018). Unfortunately, SGD has been shown to be ineffective for training transformer models, as shown in Zhang et al. (2020); Pan & Li (2023).
@@
Nevertheless, a recent study Zhao et al. (2024b) suggests a promising methodology: while SGDM generally does not work well with transformers, using SGDM for the majority of parameters and Adam for a selected subset can lead to effective training. This raises the question: Could a hybrid approach using SGD or signSGD instead of SGDM be viable? If the key subset of parameters is handled by advanced algorithms, can the other parameters be trained effectively with state-free optimizers?
@@
To address this question, we conducted an experiment on LLaMA-130M, where we utilized the Adam (Kingma, 2014) for state-full parameters and signSGD (Bernstein et al., 2018) for state-free parameters [^4] See detailed description of the setup in Section A.1. Once again we used Random projection and highlighted the result in the second part of Table 1. Full-rank updates significantly enhance performance, approaching the efficiency of the memory-intensive Adam optimizer. These findings underscore the potential of state-free algorithms for updating a substantial portion of the parameter space, paving the way for efficient and scalable optimization methods that deliver high performance without the significant memory costs traditionally associated with state-of-the-art optimizers.
4 Full-Rank Updates with GrAdient spLitting
@@
General framework. The setup outlined in the conclusion of Section 3.2 results in a general framework for memory-efficient optimization. It operates as follows: the entire space is partitioned into state-full and state-free subspaces. The state-full subspace is updated using an advanced algorithm, while the state-free subspace is updated using a state-free method. After a certain number of steps, the state-full subspace is changed to better explore the optimization space. A formal description is presented in Algorithm 1.
We note that this framework allows for variation not only in the state-full optimizer but also in the choice of projection and state-free optimizer. However, determining the optimal state-free optimizer and the projection method onto the state-full subspace is not readily apparent. In this section, we strive to find the optimal configuration.
@@
State-free optimizer. We conducted a preliminary experiment using different state-free algorithms to choose between SGD and signSGD (Bernstein et al., 2018). Table 10 shows that signSGD outperforms SGD, leading us to favor signSGD. We attribute this performance to the similarities between signSGD and Adam (Kingma, 2014), as noted in Balles & Hennig (2018); Balles et al. (2020); Kunstner et al. (2023). Additionally, signSGD produces updates of similar magnitude to those generated by Adam, which simplifies the calibration of the learning rate for state-free parameters.
@@
Projection type. When selecting a projection method, it is crucial to strike a balance between quality and memory efficiency. When using SVD decomposition for projection matrices, as in GaLore (Zhao et al., 2024a), the method better preserves the information embedded in the gradient but requires additional memory for storing projection matrices and computational resources for performing the SVD. To reduce computational demands, one could employ random coordinate projection denoted as RandK, but this requires additional memory or recomputation [^5] See Appendix C for discussion on the memory requirements for different projection methods.. A more structured alternative is to select not random entries but entire random columns or rows. The most aggressive approach follows the method from BAdam, wherein an entire block is chosen as the state-full subspace. The performance results obtained with all these variants are presented in the second part of Table 1. SVD slightly outperforms both RandK and Block projections, demonstrating comparable performance. Nonetheless, a downside is the increased compute and memory demand from SVD. Therefore, we opt for the blockwise selection, as it is the most memory-efficient — requiring only the storage of active block indices.
In experiments in Section 6, we use a specific variant with AdamW as the State-Full optimizer and signSGD as the State-Free optimizer. We primarily employ blockwise projection but switch to column-wise projection when the number of parameters in any single block exceeds memory budget, as detailed in Section 7. In addition, PyTorch-like pseudocode of our framework is presented in Appendix G.
@@
For Line 7, state projection, in Algorithm 1, we note that if the projection does not change, i.e., $P_{k,i}=P_{k-1,i}$, then $P_{k,i}(P^{-1}_{k-1,i}(s))=s$. Thus, we only need to project states when the projection changes from one round to another. However, our preliminary experiments with RandK selection showed that resetting states performs comparably to projection. Therefore, we could replace this projection with state resetting when the projection changes, which also aligns with blockwise subspace selection. However, either resetting or projecting states is important since we want projected gradients and optimizer states to reside in the same space. For instance, GaLore ignores this step, which leads to degraded performance when projections are updated frequently; see Appendix D and Section 6.4 for details.
@@
2:     $\tilde{g}^{k}\leftarrow\nabla f_{\zeta^{k}}(x^{k})$
3:     $\tilde{m}_{j}^{k}\leftarrow(1-\beta)\tilde{g}_{j}^{k}+\beta\begin{cases}\tilde{m}_{j}^{k-1}&\text{if }j\in J_{k},\\ 0&\text{otherwise;}\end{cases}$
4:     $\tilde{u}_{j}^{k}\leftarrow\begin{cases}\tilde{m}_{j}^{k}&\text{if }j\in J_{k},\\ \tilde{g}_{j}^{k}&\text{otherwise;}\end{cases}$
5:     $x^{k+1}\leftarrow x^{k}-\alpha^{k}\tilde{u}^{k}$
@@ 5 Theoretical Results
For the theoretical analysis, we consider the case where the State-Free optimizer is SGD and the State-Full optimizer is SGD with momentum (SGDM). For the projection, we use coordinate-wise projection. This special case of FRUGAL is provided in Algorithm 2. We minimize the objective
$\textstyle\min_{x\in\mathbb{R}^{d}}\big{\{}f(x):=\mathbb{E}_{\zeta{{}^{k}}}[f_{\zeta{{}^{k}}}(x)]\big{\}},$ (1)
where we access $f$ via a stochastic oracle that takes $x$ as input and returns $(f_{\zeta{{}^{k}}}(x),\nabla f_{\zeta{{}^{k}}}(x))$.
@@ 5.1 Notation and Preliminaries
$\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|.$ (2)
@@
Table 2: Comparison of validation perplexity and memory estimation for various optimization methods across LLaMA model scales trained on C4. We also indicate the additional memory overhead introduced by the optimization algorithm. The values are calculated assuming that each float value occupies 4 bytes (float32). $\rho$ denotes the proportion of the Linear layer parameters in the state-full subspace. Note that Embeddings, RMSNorms, and Output layer are always trained with AdamW.
@@
Assumption 5.1.
  1. 1.
    Smoothness: The objective $f(x)$ in equation 1 is $L$-smooth (Equation 2).
  2. 2.
  3. 3.
    Independent samples: The random samples $\{\zeta^{k}\}_{k=1}^{\infty}$ are independent.
  4. 4.
5.2 Convergence of Algorithm 2
Below, we present the main convergence theorem.
Theorem 5.2.
Let Assumption 5.1 hold and $\alpha^{k}=\alpha\leq\frac{1-\beta}{L(4-\beta+\beta^{2})}$. Then, the iterates of Algorithm 2 satisfy
$\textstyle\frac{1}{k}\sum_{i=1}^{k}\mathbb{E}[\|g^{i}\|^{2}]$ $=\mathcal{O}\biggl{(}\frac{f(x^{1})-f^{*}}{k\alpha}+$
$+L\alpha\sigma^{2}\Bigl{(}1+\frac{\hat{p}^{k}_{\max}(1-\bar{p}^{k}_{\min})\beta}{(1-\beta)}\Bigr{)}\biggr{)},$
The proof is deferred to Appendix E. Let us analyze the obtained result. Firstly, if $J_{k}=[d]$ or $J_{k}=\emptyset$, Algorithm 2 becomes SGDM and SGD, respectively. In this case, we have $\bar{p}^{k}_{\min}=1$ for SGDM and $\hat{p}^{k}_{\max}=0$ for SGD. Therefore, the resulting rate is $\mathcal{O}\left(\nicefrac{{1}}{{k\alpha}}+L\alpha\sigma^{2}\right)$, which recovers the best-known rate for both SGD and SGDM under these assumptions (Liu et al., 2020). Furthermore, if at each step each coordinate is sampled independently with probability $p$, we have $\bar{p}^{k}_{\min}=\hat{p}^{k}_{\max}=p$. Therefore, we recover the same rate if $p=\mathcal{O}\left(1-\beta\right)$ or $p=\mathcal{O}\left(\beta\right)$. Finally, in the worst case (e.g., $J_{k}$ is deterministic and $0<|J_{k}|<d$), we have $\bar{p}^{k}_{\min}=0$ and $\hat{p}^{k}_{\max}=1$. Thus, the rate becomes $\mathcal{O}\left(\nicefrac{{1}}{{k\alpha}}+\nicefrac{{L\alpha\sigma^{2}}}{{1-\beta}}\right)$, which is worse by a factor of $\nicefrac{{1}}{{1-\beta}}$. However, this is expected since the bias from momentum is not outweighed by the variance reduction effect, as only the coordinates with momentum enjoy reduced variance; see Lemmas E.2 and E.3 in the appendix for details.
6 Pre-training experiments
In this section, we evaluate the performance of FRUGAL on the language models pre-training.
6.1 Comparison to existig memory-efficient algorithms
@@
To begin, we compare our framework with existing memory-efficient methods across four sizes of LLaMA-based architectures: 60M, 130M, 350M, and 1B.
Setup. The core setup for pre-training is taken from Zhao et al. (2024a). We utilize LLaMA-based (Touvron et al., 2023a) model architectures and train them on the Colossal Clean Crawled Corpus (C4) dataset (Raffel et al., 2020). The C4 dataset is intended for pre-training, making this setup a good approximation of real-world applications. A detailed description of the setup can be found in Section A.1.
Table 3: Perplexity and memory consumption (weights, gradients and optimizer states) of different size LLaMA models pre-trained on C4 for 100k iterations (10B tokens) using AdamW with pure bf16 of mixed precision.
Model size Format Memory Perplexity
175M Mixed Precision 2.0GB 17.43
350M Pure bf16 2.1GB 17.75
350M Mixed Precision 4.2GB 15.16
1.3B Pure bf16 7.7GB 16.51
However, we made several critical modifications compared to Zhao et al. (2024a) to align the experimental setup with practical training scenarios. Below, we discuss each modification and provide a detailed rationale for these decisions.
@@
Table 4: Perplexity of LLaMA-130M models pre-trained on C4 for 100k iterations (10B tokens). The leftmost column indicates the modules moved to the state-free set and trained using signSGD. The results show that Output layer, unlike Embeddings and RMSNorms, are exceptionally responsive to the choice of optimization algorithm from AdamW to signSGD.
@@
Linear, Output layer 34.66
  • Training Duration. The training approach in Zhao et al. (2024a) aligns with the empirical rule from scaling laws (Hoffmann et al., 2022), which suggests using approximately 20 times the size of the model in tokens for training. However, this number of tokens is far from achieving convergence. In practice, models are typically trained for significantly longer periods (Touvron et al., 2023b; Zhang et al., 2024b). One reason for this discrepancy is that the original scaling laws do not account for the inference of the model after training (Sardana & Frankle, 2023). For our experiments, we chose 200k steps for the 60M and 130M models, 240k for the 350M model, and 300k for the 1B and 3B models.
  • Mixed Precision. Pure 16-bit training has been shown to potentially compromise model convergence and accuracy (Zamirai et al., 2020). This degradation occurs because formats such as float16 or bfloat16, used to store master weights, lack the numerical precision needed for accurate and fine-grained weight updates. Consequently, mixed precision training has become a more common approach for training language models (Le Scao et al., 2023; Almazrouei et al., 2023). Moreover, even when training with fp8, the master weights are typically stored in fp32 format (Liu et al., 2024a).
    Our experimental results strongly support the importance of precision choice: in Table 3 we show that adopting pure bf16 training led to such significant performance degradation that doubling the model size failed to compensate for it, effectively negating any memory benefits from reduced precision storage. While training in pure 16-bit format is also possible, stochastic rounding (Gupta et al., 2015; Zamirai et al., 2020) is often employed to mitigate the aforementioned issue. Given that the goal of this research is to identify the optimal optimization algorithm, we deemed it more appropriate to compare optimizers in a transparent and stable setup that does not require auxiliary tricks. Hence, we primarily used Mixed Precision training for its illustrative value in understanding each method’s potential. However, for completeness, we also conducted experiments in pure bfloat16 format, detailed in our ablation study Section 6.4.
Baselines. We use the following methods as baselines:
Table 5: Pre-training LLaMA 3B on C4 dataset for 300K steps. Validation perplexity for different iterations is reported.
Method 100k 200k 300k
AdamW 14.2 12.25 10.93
FRUGAL, $\rho=0.25$ 14.33 12.42 11.07
FRUGAL, $\rho=0.0$ 14.78 12.76 11.35